- 2/28-3/15: Health check and maintenance for VIS
PRM: One flag was dropped. -> leave it
3/4-14 PRM: Misalignment does not reach where beam hits damper.
3/6 PR3: range -> incline by 20-30mrad at mirror because of bistable.
3/15: BS: Yaw offset adjustment by hand. -> 3/5-6
3/12: BS ok
3/15: some SRs work
3/5 afternoon or 3/6 morning: Coolers for ITMs will be stopped for 2-3 hours to change water.
3/22: Pump down
- 3/27: Fixing PR2
- 4/3: Closing and pump down
3/19: Cooling stop
3/25: Venting start
- 3/27: Opning chamber
- 3/27~: Fixing thermal touch
- ~4/3: closing chamber and start pumping down.
- 4/12: Cooling starts.
2/11-2/14: Xend NAB installation
2/21: Removing FC, Payload release
2/27: ETMX release
2/23-3/6: Xend Pcal finalization 1
3/6: Beam position by IR and Green
3/5 AOS: EY and EX(if they have time) work.
3/6: Green X alignment (no IR), and no hazard at Y end.
3/11, ~20: Xend WAB finalization.
3/11~3/15: ETMX tuning.
3/19: ETMX finalization by Takahashi
3/18,19: ETMX alignement
3/23-24: Pcal alignment
3/24: Removing PD
3/25: closing and pump down
- 4/12: Cooling start
1/25-2/20: Yend Pcal finalization
3/5: EYA remaining works by Enomoto, Inoue.
2/25-3/5: Yend WAB finalization
3/6: Y end closing (no laser hazard at least in the morning)
3/7: Y end pump down
3/22-23: leak test
3/26: Venting start
- 3/28~: Fixing short inside chamber
- ~4/2: closing chamber and start pumping down.
- 4/10: Cooling starts.
- 4/1~ : Night work starts, daily meeting will be moved to 1PM
- Summary of work will last for a while in the morning meeting.
2/18-2/22: PRMI commissioning
- End of April~: DRMI commissioning
- End of April~: Y arm commissioning
- ~End of May: FPMI commissioning
- Official night work starts 4/1.
Report of work
VAC:ETMY venting which was stoped the end of yesterday was restarted this morning. -> needs monitor (Lucia).
- DGS: HDD on k1grd0 was done this morning. Guardian is back now.
- CRY: ETMY chamber will be opened today and main work will be performed tomorrow.
- CRY: ITMX will be investigated to see the cause of not going low temperature.
- Type Bp: PR2 will be investigated to fix hitting somewhere.
- IOO: High power test will be performed from the end of today or tomorrow.
- DGS: Matlab license on WSs will be updated today.